OpenStudy (anonymous):

Find the sum of pentagon’s angle measures. 180 degrees 360 degrees 540 degrees 720 degrees

OpenStudy (anonymous):

@MARC_

OpenStudy (hockeychick23):

540

OpenStudy (hockeychick23):

you can split up a pentagon into 3 triangles to make it easier if you wanted to, triangles are 180 degrees so you do 180*3= 540

OpenStudy (anonymous):

thnx

OpenStudy (hockeychick23):

no problem :)

OpenStudy (cwrw238):

the general formula for sum of interior angles of a polygon is S = 180(n - 2) so for a pentagon n = 5 S = 180(5 - 2)